Experiments with rats show that if rats are dosed with 3.0 mg/kg of cocaine ( that is, 3.0 mg cocaine per kg of animal mass), the concentration of dopamine in their brains increases by 0.75LaTeX: \muμM after 60 seconds. Calculate how many molecules of dopamine would be produced in a rat ( average brain volume 5.00 mm3) after 60 seconds of a 3.0 mg/kg dose of cocaine.
Answer #1

Volume of brain in mL = Volume in mm3 * 10-3 = 5 * 10-3 mL

Increase in moles of dopamine = molarity increase * Volume in Liters = 0.75 * 10-6 M * 5 * 10-3 * 10-3 L = 3.75 * 10-9 moles

Number of molecules = avogadro number * Moles = 6.022 * 1023 * 3.75 * 10-9 = 2.26 * 1015 molecules